OpenStudy (anonymous):

find an equation of the line containing the given points. Express your answer in standard form. (2,-7) and (9,-49)

OpenStudy (anonymous):

Do I need put this equation Y-Y1= m (X-X1)

OpenStudy (lgbasallote):

that is for a point and a slope given. SInce you're given two points here, you use two-point form \[\huge y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x-x_1)\] got it?

OpenStudy (anonymous):

Find slope m = Δy/Δx first then plug m and the coordinate value !

OpenStudy (anonymous):

The answer will be 6x+y+5 right?

OpenStudy (lgbasallote):

hmm that doesn't sound right

OpenStudy (anonymous):

Why don't you just follow the step, showing your work?

OpenStudy (lgbasallote):

remember: x1 = 2 y1 = -7 x2 = 9 y2 = -49 plug them in \[\Large y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x-x_1)\]

OpenStudy (anonymous):

Y+7=-49+7/7 (X-2)

OpenStudy (anonymous):

Y=-6(X-2)

OpenStudy (anonymous):

6X+Y+5

OpenStudy (lgbasallote):

you made a sign error there @sabasyed

OpenStudy (anonymous):

y + 7 = -6 ( x -2)

OpenStudy (anonymous):

6x+Y-5

OpenStudy (lgbasallote):

better

OpenStudy (lgbasallote):

don't forget equal sign

OpenStudy (anonymous):

y = -6x +12 - 7 = -6x +5